OpenStudy (anonymous):

Andrew has two part time jobs. One pays $6 per hour and the other pays $10 per hour. He wants to make $366 per week. Write an equation in standard form that describes this situation. If he is only allowed to work 15 hours per week at the $10 per hour job, how many hours does he need to work per week in his $6 per hour job in order to achieve his goal? How do you write out the answer cause i don"t understand hoe everything would be placed??

OpenStudy (henrietepurina):

6x + 10 * 15 = 366

OpenStudy (anonymous):

someone said that there would be a lot of wording to go with it, or is it just the problem?

OpenStudy (henrietepurina):

And the answer is 36 BTW

OpenStudy (henrietepurina):

I dont think its actually that hard!

OpenStudy (anonymous):

thank you

OpenStudy (henrietepurina):

You welcome

OpenStudy (mathmale):

@msumner: You have a math problem involving TWO variables here. Why not call them x and y? The amount Andrew earns from the $6/hour job is \[\frac{ $6 }{ 1~hr }x.\]If y represents the mount he earns from the other job, write a similar expression in y instead of in x.

OpenStudy (anonymous):

would 36 = x

OpenStudy (mathmale):

@msummer: Please, if you want to understand this math problem, answer my question. I don't think it would help you if I were to give you merely "yes" or "no" as an answer.

OpenStudy (anonymous):

im sorry im trying to understand

OpenStudy (mathmale):

Andrew's other job pays $10/hour. So, his earnings at that job are\[\frac{ $10 }{1~hr }y\]where y is the number of hours he works there. Add these two amounts together, please, and equate them to the $366 Andrew wants to earn.

OpenStudy (acxbox22):

this does not require 2 variables multiply 15 by 10=150 15 is # of hours at $10 job subtract 150 from 366 which equals 216 divide 216 by $6 job to find how many hours you work there to meet your goal just like @HenrietePurina

OpenStudy (anonymous):

oh that makes sence so it =36

OpenStudy (mathmale):

@acxbox22: Certainly that is true, but the approach I've advocated is more typical of more advanced algebra problems in which there are constraints to be satisfied. Would you mind waiting until @msumner and I are finished with this approach before discussing yours? I know you were trying to be helpful, but you have interrupted a discussion in midstream.

OpenStudy (anonymous):

I mean no disrespect but I understood @acxbox22 more then you, its not that you didn't explain it right it just he made more sense to me

OpenStudy (acxbox22):

i was just trying to give msumner an easier approach, one which she can understand

OpenStudy (anonymous):

thank you both for your help

OpenStudy (mathmale):

Fine, so be it. Let's not interrupt one another next time, all right?

OpenStudy (acxbox22):

as you wish Moderator